The sum of two brothers is 46. The younger brother is 10 more than a third of the older brothers age. How old is the younger brother?
Answer by 1335(25)

Call the older brother x and then write the second brother's age in terms of the first.
x+1/3x+10=46
Simplify: 4/3x+10=46
4/3x=36
x=27 -the older brother.
1/3(27)+10=9+10= 19 -the younger brother